OpenStudy (anonymous):

240yards/1hour find the unit rate in feet per second?

OpenStudy (anonymous):

1 yard = 3 feet 1 hour= 60 minutes = 3600 seconds

OpenStudy (anonymous):

can you solve it now?

OpenStudy (anonymous):

@vikayefremova

OpenStudy (anonymous):

well...if there's 3600 seconds in an hour then.. 3600/240 yrds = 15 i think that's the answer :/

OpenStudy (anonymous):

i got 0.2 feet

OpenStudy (anonymous):

the answer should be ft/s @vikayefremova

OpenStudy (anonymous):

i'm not sure sorry :/

OpenStudy (anonymous):

@pudding got the same answer

OpenStudy (anonymous):

ysy

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

@vikayefremova \[\frac{240 yd}{1 hr}*\frac{3 ft}{1 yd}*\frac{1hr}{3600\sec}=.2\frac{ft}{\sec}\]

OpenStudy (anonymous):

??? i got \[\frac{ 720 }{ 3600 }\]

OpenStudy (anonymous):

ohhh, that's how you do it! i wasn't really sure...well there's your answer @pudding

OpenStudy (anonymous):

thank you xD

OpenStudy (anonymous):

its the same thing pudding

OpenStudy (anonymous):

i see o.o u did it another way but we got the same answer @completeidiot

OpenStudy (anonymous):

if you ever take chemistry, you will probably end up writing like that it helps

OpenStudy (anonymous):

thnks m in Algebra 1 right now though what grade r u in?

OpenStudy (anonymous):

3rd yr engineering student

OpenStudy (anonymous):

nice it's my 2st time using this site and i love it! thnks for the help